# GARCH copulas and GARCH-mimicking copulas

### v1

@@ 5.1 Mimicking jointly symmetric GARCH copulas
    the copula of a spherically symmetric distribution;
@@
    an equal (25/25/25/25) mixture of any copula and its rotations through 90, 180 and 270 degrees;
  4. 4.
    a copula constructed from another copula $C^{*}$ by inverse v-transformation with the symmetric v-transform, i.e. a copula with density of the form $c^{*}(|2u-1|,|2v-1|)$.
In view of Proposition 6, the densities of the copulas arising from the first 3 approaches can also be written in the form $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some implicit copula $C^{*}$ with density $c^{*}$, so the 4th approach should be considered as the overarching approach. Proposition 3, part (ii), and the lower-left panel of Figure 3 give clues about the kind of copula density $c^{*}$ we should be targeting. In particular, we note that there should be strong dependence in the upper right tail and some asymmetry.
An example of (1) is the copula of a spherical t distribution, i.e. a bivariate t distribution located at the origin with identical marginal distributions and correlation equal to zero. We call this copula the spherical t copula and denote it $C^{t}_{\nu}$. Since a spherical t-distributed random vector $(Y,Z)^{\top}$ is jointly symmetric, so is its copula. As the degree-of-freedom parameter tends to zero, this copula becomes increasingly cross-shaped and converges to an equal mixture of the comonotonicity and countermonotonicity copulas (McNeil et al.,, 2022).
It is of interest to identify the copula corresponding to v-transformation of $C^{t}_{\nu}$, in other words, the copula of $(|Y|,|Z|)^{\top}$. We call this copula the absolute spherical t copula (ast) and denote it $C^{|t|}_{\nu}$. By Corollary 1 and (16) its density must satisfy
@@
Integrating any one of these equalities and observing the boundary conditions for a copula gives us an expression for $C^{|t|}_{\nu}$. For example, integrating the final equality gives
$C^{|t|}_{\nu}=4C_{\nu}^{t}\left(\frac{1+u}{2},\frac{1+v}{2}\right)-u-v-1.$ (20)
This one-parameter copula, which appears to be new, interpolates between independence (as $\nu\to\infty$) and comonotonicity (as $\nu\to 0$). For small degrees of freedom, the density $c^{|t|}_{\nu}$ does bear some resemblance to the lower-left panel of Figure 3, but cannot reproduce the non-exchangeability.
An example of (2) would be an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and the t copula $C^{t}_{\nu,-\rho}$ for some $\rho\geqslant 0$; obviously this contains the spherical t copula $C^{t}_{\nu}$ as the special case where $\rho=0$. Examples of (3) could be constructed using the Gumbel, Clayton or Joe copulas, to take advantage of the strong tail dependence of these copulas in either the upper-right or lower-left corners; this includes the equally-weighted special case of the so-called convex Gumbel copula (Junker and May,, 2005) used by Loaiza-Maya et al., (2018) for financial return data. Mixtures of types (2) and (3) all imply v-transformed copulas $C^{*}$, whose forms may be easily derived using the method we have used for the spherical t copula. If we follow the explicit construction (4), we need to choose a copula $C^{*}$ that has stronger dependence in the upper tail than the lower tail. This suggests considering the Gumbel, Joe or 180-degree-rotated (survival) Clayton copulas.
So far none of these constructions address the issue of non-exchangeability which was clearly evident in certain examples in Figure 2. One possible construction for non-exchangeable copulas is the one proposed by Khoudraji, (1995) and later generalized by Liebscher, (2008). In this construction two copulas $C_{A}$ and $C_{B}$ are combined to give a new copula
$C_{K}(u,v)=C_{A}(u^{1-a_{1}},v^{1-a_{2}})C_{B}(u^{a_{1}},v^{a_{2}})$ (21)
by choosing constants $a_{i}\in[0,1]$ for $i=1,2$ and the resulting copula $C_{K}$ is non-exchangeable when $a_{1}\neq a_{2}$. In our use of this construction we set $C_{B}$ to be the independence copula. We apply the Khoudraji extension to copulas before either mixing them with their rotations or applying inverse v-transformation.
We note, however, that the class of copulas in (21) does have one problematic feature for maximum-likelihood inference; even when $C_{B}$ is the independence copula, in order to evaluate the copula density of $c_{K}$ of $C_{K}$, we need to be able to evaluate, not only the copula density $c_{A}$, but also the copula $C_{A}$ and its partial derivatives. This is a particular challenge for copulas related to the t copula [^3] For example, existing R software for evaluating the distribution function of the t copula imposes some limitations on $\nu$. The algorithms in the R package copula (Hofert et al.,, 2023), which are based on methods by Gentz and Bretz (Gentz and Bretz,, 2009), are limited to integer $\nu$ and this is insufficiently fine-grained for applications of $C^{|t|}_{\nu}$. The C++ library underlying rvinecopulib (Nagler and Vatter,, 2022) imposes the restriction that $\nu\geqslant 2$..
5.2 Simulation study of jointly symmetric case
We explore the ideas of the previous section via a Monte Carlo simulation study. We generate data from the classical symmetric ARCH(1) and GARCH(1,1) processes using the parameter values and innovation distributions summarized in Table 1. We then fit copulas with the property of joint symmetry to the generated realizations of the pairs $(X_{t},X_{t+1})^{\top}$.
In this and all subsequent simulations, the analyses are based on simulated samples of 50000 points. Copulas are estimated by the method of pseudo maximum likelihood (PML) (Genest and Rivest,, 1993); in other words, marginal distributions are estimated empirically, the data are mapped to the unit square using the empirical distribution function and copula parameters are then determined by maximum likelihood. Models are compared according to their AIC values.
The first copula fitted in Table 2 is $C^{t}_{\nu}$, the copula of the spherical t distribution, i.e. the t copula with the correlation parameter fixed at $\rho=0$; the only free parameter is the degree-of-freedom parameter $\nu$. We also label this copula iv-ast, to emphasize that it can be viewed as the inverse v-transformed version of the absolute spherical t copula $C^{|t|}_{\nu}$. The second model is the generalisation of $C^{t}_{\nu}$ consisting of an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and its 90-degree rotation $C^{t}_{\nu,-\rho}$ for $\rho\geqslant 0$.
According to AIC, the $C_{\nu}^{t}$ is the best model in the table in two of six cases, which are the cases where the innovations are Student t4. The mixture of two t copulas is an inferior model to $C_{\nu}^{t}$ in 5 of 6 cases, the only exception being GARCH with normal innovations. The maximum likelihood method for the t mixture returns an estimate $\rho\approx 0$ in those 5 cases, which accounts for the fact that the difference in AIC values is approximately 2.
The next panel of 3 models contains equal mixtures of the Clayton, Gumbel and Joe copulas and their 90, 180 and 270 degree rotations. The Gumbel case is the equally-weighted special case of the convex Gumbel copula of Junker and May, (2005), as also used by Loaiza-Maya et al., (2018). The next 3 copulas in the first panel of Table 2 are inverse-v-transformed versions of the upper tail-dependent Gumbel, Joe and Clayton survival copulas (the latter described as Clayton-180), i.e. copulas with densities of the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of either the Clayton survival copula, the Gumbel copula or the Joe copulas.
Comparing all of the 1-parameter models, the $C_{\nu}^{t}$ (iv-ast) copula is always the best model; second place is usually occupied by the iv-Joe copula, except for ARCH with Gaussian innovations where the iv-Clayton180 copula is favoured. Interestingly, while the iv-Gumbel copula is the worst of the inverse-v-transformed copulas, the mixture of Gumbels is often the best of the mixtures.
The remaining models in Table 2 are non-exchangeable. The fourth panel of the table contains mixture copulas, except that the Khoudraji extension in (21) is applied to each copula before the mixture of the 4 different rotations is constructed. Thus these models introduce two further parameters $a_{1}$ and $a_{2}$. The final panel contains inverse-v-transformed versions of the Khoudraji extensions of the Clayton survival, Gumbel and Joe copulas. It would be appealing to add the Khoudraji construction based on the iv-ast copula, but this model is computationally more demanding and not currently a viable candidate model for practical applications.
The best of the 3-parameter non-exchangeable models are always in the inverse-v-transformed class. The mixture copulas are only really competitive for t2.5 distributed innovations. The best-fitting copulas for Gaussian innovations (iv-K-Clayton-190 for ARCH and iv-K-Joe for GARCH) give estimates of $a_{2}$ which are very close to zero, while the best-fitting copula for t2.5 innovations (iv-K-Joe) give estimates of $a_{1}$ which are very close to zero. This makes sense in relation to Figure 2 since the non-exchangeability in the copulas seems to be caused by a “twist” in different directions.
It is notable that the best model of all is iv-ast in the two cases where the innovations are t4. In the case of ARCH, this is in accord with the near-exchangeability evident in Figures 1 and 2. The estimated degree of freedom parameters are 2.56 for ARCH and 2.67 for GARCH.

### v2

@@ 4.1 Proposal for a simplified D-vine structure
    We choose pair copula densities of the form $c_{k}(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2});\theta_{k})$, based on two linear v-transforms and a copula family $C^{*}(u,v;\theta)$ satisfying $c^{*}(0,0;\theta)<\infty$ which can model the full range of positive dependence between independence and comonotonicity for different values of the parameter $\theta$ in a parameter space $\Theta$; examples are the ast, Joe and survival Clayton copulas. This construction offers flexibility at the cost of only three parameters: $\theta_{k}$ controls the strength of dependence while $\delta_{1}$ and $\delta_{2}$ accommodate possible asymmetries.
@@
    We model the parameters $(\theta_{k})_{k\in\mathbb{N}}$ as a parametric function of the lag $k$ so that $\theta_{k}=g(k,\bm{\vartheta})$ where $\bm{\vartheta}$ is a low-dimensional parameter vector. More details of this parameterisation are given below.
@@
    We assume that the asymmetry parameters $\delta_{1}$ and $\delta_{2}$ are constant for all lags $k$. This strong assumption is based on the observation that these parameters are equal to 0.5 at all lags in jointly symmetric models. It is also partly influenced by the GARCH(1,1) model where a single parameter $\gamma_{1}$ controls leverage and a further parameter $\lambda$ controls asymmetry of innovations.
There is another practical reason for choosing linear v-transforms. After making the simplifying assumption, the copula density implied by the D-vine in equation (1) takes the form
$c_{U_{1},\ldots,U_{n}}(u_{1},\ldots,u_{n})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{U_{1}|\bm{U}_{[2:k]}}(u_{j-k}\mid\bm{u}_{[j-k+1:j-1]}),F_{U_{k+1}|\bm{U}_{[2:k]}}(u_{j}\mid\bm{u}_{[j-k+1:j-1]})).$ (18)
It is well known that the the evaluation of the conditional distribution functions $F_{U_{1}|\bm{U}_{[2:k]}}$ and $F_{U_{k+1}|\bm{U}_{[2:k]}}$ in (18) is straightforward if we have a method of computing the so-called $h$-functions of the copulas $C_{k}$, which are the partial derivatives $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$; see, for example, Aas et al., (2009) or Smith et al., (2010). This is easy in the linear case but, in the non-linear case, cumbersome numerical integration seems in general to be necessary. More details of the formulas for the linear case are found in Section S6.1 of the Supplementary Material while in Section S6.2 we show that, in the special case of two identical linear v-transforms, D-vines with inverse-v-transformed copulas also belong to the family of vt-D-vines proposed in Bladt and McNeil, 2022b .
The relationship $\theta_{k}=g(k,\bm{\vartheta})$ between the parameters $\theta_{k}$ and the lag $k$ is determined implicitly using a method developed in Bladt and McNeil, 2022a and Bladt et al., (2025). Let $\tau^{*}:\Theta\to[0,1)$ be the invertible function that maps the parameter value of the copula $C^{*}(u,v;\theta)$ into the Kendall’s tau measure of association and write $(\tau^{*})^{-1}$ for its inverse. Let $(\omega_{k}(\bm{\vartheta}))_{k\in\mathbb{N}}$ denote the partial autocorrelation (pacf) of a classical Gaussian ARMA($p$,$q$) process, where $\bm{\vartheta}$ denotes the vector of $p+q$ parameters and where we assume that the parameters are constrained to ensure $\omega_{k}(\bm{\vartheta})\in[0,1)$ for all $k\in\mathbb{N}$. We determine the sequence of parameters $(\theta_{k})_{k\in\mathbb{N}}$ by setting $\theta_{k}=(\tau^{*})^{-1}\left(2\pi^{-1}\arcsin(\omega_{k}(\bm{\vartheta}))\right)$ for all $k$. This means that we equate the Kendall’s tau values of the copula sequence $(C^{*}(u,v;\theta_{k}))_{k\in\mathbb{N}}$ with the Kendall’s tau values of the sequence of pair copulas that characterise the Gaussian ARMA($p$,$q$) model. In practice we favour the ARMA(1,1) model based on the fact that the autocorrelation function of the squared values of the classical GARCH(1,1) process decays like an ARMA(1,1) process, as noted in Section 2.4.3. We also consider AR($p$) models to faciliate comparison with D-vine($p$) models in the literature; these lead to the estimation of $p$ parameters $\theta_{1},\ldots,\theta_{p}$.
4.2 Examples
To illustrate the potential of the D-vine models proposed in Section 4.1, we provide four examples, three using simulated data from the GARCH(1,1)-type processes documented in Table 1 and one using real daily return data for the USD/AUD exchange rate, as previously analysed in Loaiza-Maya et al., (2018). The simulated datasets consist of 10000 values from the GARCH(1,1) model with (i) t4 innovations, (ii) skewed t4 innovations and (iii) skewed t4 innovations and leverage. The exchange-rate returns consist of 3669 values from 2 January 2001 to 7 August 2015.
Models are fitted using the generalization of the two-stage method of Chen and Fan, (2006) to higher-order D-vines. For the simulated data, the marginal distribution function is estimated by the empirical distribution function while, for the real data, we use the same adaptive kernel density estimator as Loaiza-Maya et al., (2018) to facilitate comparison of results. The maximum-likelihood fit of the parametric copula model to the pseudo-copula observations can be accomplished in a single step, rather than by using the common stepwise procedure where the sequence of $C_{1},\ldots,C_{p}$ defining a D-vine($p$) model is estimated in $p$ successive steps.
To all datasets we fit D-vine(1) and D-vine(5) models based on t copulas (T1, T5) and convex Gumbel mixtures (B1, B5), as used in Loaiza-Maya et al., (2018). For the new models of Section 4.1, we use inverse-v-transformed Joe, survival Clayton and ast copulas and parameterizations based on the partial acfs of Gaussian AR(1), AR(5) and ARMA(1,1) processes. The copula sequence in the ARMA(1,1) case is truncated at 40 for the real data and 20 for the simulated data, since persistence of volatility is weaker for the latter. For the empirical data only, we add D-vine(1) and D-vine(5) models based on mixtures of two t copulas (A1, A5) to facilitate further comparison with models in Loaiza-Maya et al., (2018). For the simulated data, we add a D-vine(2) model with two t copulas, as suggested by Zhao et al., (2022) for GARCH approximation. Since the different models have very different numbers of parameters, we compare models using both AIC and BIC, where the latter makes greater adjustments for overfitting.
For all four datasets the best model, as selected by both the AIC and BIC information criteria, is ast-ARMA(1,1) models and we show estimated parameter values and standard errors in Table 3. It may be noted that the estimates of $\delta_{1}$ and $\delta_{2}$ are both close to 0.5 for the symmetric GARCH model, but more offset in the other cases. The estimates of the AR and MA parameters ($\phi$ and $\psi$) are larger in absolute value for the real data than for the GARCH simulations, implying a much slower decay of the partial acf and greater persistence in volatility. We also give the parameters $\nu_{1}$ and $\nu_{2}$ of the ast copulas at the first two lags as well as the corresponding values of Kendall’s tau, $\tau_{1}$ and $\tau_{2}$.
Dataset $\phi$ $\psi$ $\delta_{1}$ $\delta_{2}$ $\nu_{1}$ $\nu_{2}$ $\tau_{1}$ $\tau_{2}$
GARCH(1,1), t4 innovations 0.780 -0.614 0.510 0.515 2.96 5.07 0.132 0.079
(0.012) (0.016) (0.017) (0.016)
GARCH(1,1), skew t4 innovations 0.786 -0.614 0.493 0.427 2.82 4.85 0.139 0.082
(0.012) (0.017) (0.016) (0.016)
GARCH(1,1), skew t4 + leverage 0.790 -0.607 0.644 0.434 2.60 4.56 0.150 0.088
(0.013) (0.018) (0.015) (0.014)
USD/AUD daily exchange-rate returns 0.982 -0.934 0.528 0.446 5.82 6.59 0.069 0.061
(0.0029) (0.0073) (0.024) (0.023)
Table 3: Parameter estimates and (standard errors) for D-vines based on ast copula and partial acf of ARMA(1,1) model fitted to four datasets. Models have four primary parameters, $\phi$ and $\psi$ from the ARMA(1,1) partial acf and $\delta_{1}$ and $\delta_{2}$ from the linear v-transforms; $\nu_{1}$ and $\nu_{2}$ are implied parameters of the ast copulas at the first two lags and $\tau_{1}$ and $\tau_{2}$ are the corresponding values of Kendall’s tau.
The full model comparisons for the USD/AUD data are found in Table 4 and we give more commentary on these results since they are broadly representative of all four examples; corresponding tables for the three simulated datasets are given in Section S7 of the Supplementary Material. Comparing D-vine(1) models, the copulas based on inverse-v-transformation of the Joe and ast copulas lead to higher log-likelihoods and lower AIC and BIC values than the convex Gumbel mixture model B1. This suggests the fundamental shape of these copulas is slightly better than the mixture copulas for modelling $C_{1}$. Comparing D-vine(5) models, B5 gives the highest log-likelihood value but is a very expensive model to fit in terms of the number of parameters and the time required to maximize the likelihood [^3] Loaiza-Maya et al., (2018) fit this model using MCMC in the Bayesian paradigm and we use their posterior mean estimates as starting values in maximum likelihood estimation.. The D-vine(5) models based on inverse-v-transformed copulas have less than a third of the parameters, are quicker to fit and lead to lower values for the information criteria. However, our favoured D-vine(40) models based on the partial acf of an ARMA(1,1) model lead to the largest log-likelihood values and the lowest AIC and BIC values at the cost of only 4 parameters.
Loaiza-Maya et al., (2018) show that their favoured models (B5 and T5) lead to one-step-ahead forecasts of conditional quantiles (value-at-risk) that are more accurate than the ones produced by classical symmetric GARCH models and are not rejected by Christoffersen’s test of conditional coverage at the 99% level (Christoffersen,, 1998). This is also true of our favoured models (Joe-AR(5), ast-AR(5), Clayton180-ARMA(1,1) and ast-ARMA(1,1)); results are documented in Section S8 of the Supplementary Material.
Model D-vine order No. pars Loglik AIC BIC
T1 1 2 36.72 -69.43 -57.01
A1 1 5 36.71 -63.42 -32.38
B1 1 5 38.70 -67.40 -36.36
Joe-AR(1) 1 3 41.17 -76.33 -57.71
Clayton180-AR(1) 1 3 38.26 -70.51 -51.89
ast-AR(1) 1 3 41.13 -76.27 -57.64
T5 5 10 153.04 -286.08 -224.01
A5 5 25 155.63 -261.26 -106.07
B5 5 25 163.09 -276.17 -120.98
Joe-AR(5) 5 7 154.13 -294.26 -250.81
Clayton180-AR(5) 5 7 151.32 -288.63 -245.18
ast-AR(5) 5 7 155.72 -297.45 -253.99
Joe-ARMA(1,1) 40 4 247.21 -486.42 -461.59
Clayton180-ARMA(1,1) 40 4 255.29 -502.59 -477.76
ast-ARMA(1,1) 40 4 261.91 -515.83 -491.00
Table 4: Comparison of D-vine models in Loaiza-Maya et al., (2018) and the D-vine models of Section 4 fitted to return data for the USD/AUD exchange rate for the period 2 January 2001 to 7 August 2015; the marginal distribution is estimated nonparametrically using adaptive kernel density estimation. T1, A1 and B1 are D-vine(1) models based, respectively, on t copulas, mixed t copulas and mixed convex Gumbel copulas as used in Loaiza-Maya et al., (2018), while T5, A5 and B5 are D-vine(5) extensions. The models following the construction in Section 4.1 are described by the name of the copula $C^{*}$ and the name of the AR or ARMA model underlying the parameterisation $\theta_{k}=g(\bm{\vartheta})$; the copula sequence in the ARMA(1,1) case is truncated at order 40.
5 Conclusion
This paper has asked three questions. First, what are the bivariate copulas that underlie the serial dependencies and conditional dependencies of pairs of variables $(X_{t},X_{t+k})$ in first-order GARCH(1,1) models? Second, since the true pair copulas and conditional copulas of GARCH-type processes are inaccessible for practical use, can they be approximated by simpler constructions? Third, can these ideas help us to construct simplified D-vines that behave like GARCH-type models, but offer more flexible models for real data showing stochastic volatility?
With regard to the first question we find a number of symmetries in the dependence structures of the models with symmetric innovation distributions, even when leverage effects are introduced, but these symmetries disappear in the presence of asymmetric innovations, which are often necessary to model real-world data in practice. What can always be said is that the underlying bivariate copulas are cross-shaped, in the sense that there is a tendency for points to accumulate in all fours corners of the unit square, and that this tendency increases at shorter lags as the size of the ARCH effect increases and persists at longer lags as the aggregate size of the ARCH and GARCH effects increases.
Motivated by the recognition that v-transforms are implicitly embedded in the copulas of symmetric GARCH models, we propose a construction to approximate GARCH pair copulas based on v-transforms and positive-dependence copulas with strong dependence in the upper joint tail. Moreover, we identify a new copula - the absolute spherical t copula - which offers a promising basis for the construction.
With regard to the final question, we believe that the simplified D-vine models proposed in Section 4.1 offer effective descriptions of GARCH-like behaviour and have potential for modelling volatile financial data. Their advantages come not only from the particular choice of pair copulas, but also from a method of parameterization, using the partial acf of Gaussian ARMA processes, which gives models with a similar degree of parsimony to GARCH processes. In combination with the flexibility to choose an appropriate marginal distribution, these models could be extremely effective when data depart from the power-lawed marginal behaviour imposed by a GARCH process, as discussed in Bladt and McNeil, 2022b . However, the proposed models also need more extensive empirical testing. It is important to conduct experiments with a wider variety of financial return data, such as stock-price, index, commodity-price and volatility-index returns. This is the subject of future work.
Another issue requiring further research is the construction of D-vines to model ARMA-GARCH behaviour. Real-world financial return series sometimes show serial correlation in the raw data as well as in the squared or absolute value series and this is usually modelled by combining ARMA and GARCH effects. In Section S9 of the Supplementary Material we consider an AR-ARCH model and suggest that mixtures of standard positive-dependence copulas and the pair copulas of this paper could be successful in addressing combined ARMA and GARCH effects.
Supplementary Material
Appendix S1 Correlation and rank correlation for h-symmetric distributions
Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined). Perhaps surprisingly, h-symmetry on its own is a sufficient condition for an absence of both correlation and rank correlation; the same is true of v-symmetry.
Proposition S1.
Proof.
We give the proof in the case of h-symmetry, the method for v-symmetry being identical. If $(Y,Z)^{\top}$ is h-symmetric about $b$ then $\mathbb{E}(Z)=b$ and we can write $\operatorname{cov}(Y,Z)=\mathbb{E}(YZ)-\mathbb{E}(Y)\mathbb{E}(Z)=\mathbb{E}(Y(Z-b))+\mathbb{E}(Yb)-\mathbb{E}(Y)b=\mathbb{E}(Y(Z-b))$. However, h-symmetry implies $\mathbb{E}(Y(Z-b))=\mathbb{E}(Y(b-Z))=-\mathbb{E}(Y(Z-b))$ implying that $\mathbb{E}(Y(Z-b))=0$ from which the result follows. ∎
Proposition S2.
Proof.
Again it suffices to give the proof in the case of h-symmetry. Let the unique copula of $(Y,Z)^{\top}$ be denoted $C$. For Spearman’s rank correlation we need to calculate $\rho_{S}(Y,Z)=12\int_{0}^{1}\int_{0}^{1}(C(u,v)-uv)\mathrm{d}u\mathrm{d}v$. From the identity C(u,v)=uC(u,1v)C(u,v)=u-C(u,1-v) we infer that
where we have made the change of variable y=1vy=1-v. We must have $\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v=\tfrac{1}{4}$ and then simple calculations give $\rho_{S}(Y,Z)=0$.
In the case of Kendall’s rank correlation we need to compute $\tau(Y,Z)=4\mathbb{E}(C(U,V))-1$ where $(U,V)^{\top}$ is a pair of variables with distribution function $C$. The identity C(u,v)=uC(u,1v)C(u,v)=u-C(u,1-v) implies that $(U,V)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,1-V)^{\top}$; indeed $(U,V-0.5)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,0.5-V)^{\top}$ since $(U,V)^{\top}$ is h-symmetric about 0.5. Hence we infer that
which implies that $\mathbb{E}(C(U,V))=\tfrac{1}{4}$ and $\tau(Y,Z)=0$. ∎
Appendix S2 More illustrations of joint densities and copulas
Figure S1: Contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values. Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations. Bottom right: classical GARCH(1,1) process with Gaussian innovations.
Figure S1 shows contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for a number of symmetric GARCH(1,1)-type processes as described in Sections 2.4.12.4.3; these pictures are counterparts to the copula pictures in Figure 1 and show the joint symmetry of the distributions. It is notable that the contours in the top right picture for an ARCH(1) model with Student t4 innovations look almost circular, showing the distribution is very close to spherical; the corresponding copula density plot in Figure 1 suggests the copula is almost exchangeable.
In the left panel of Figure S2 we show a contour plot of the absolute value copula (the copula of $(|X_{1}|,|X_{2}|)^{\top}$) for the ARCH(1) model with Gaussian innovations described in Section 2.4.1. This can be thought of as a blown-up version of the top right quadrant of the copula density for $(X_{1},X_{2})^{\top}$ and clearly shows non-exchangeability and stronger dependence in the upper tail. Note that this picture also illustrates the density of the volatility copula of ARCH(1), that is the copula of $(\sigma_{1},\sigma_{2})^{\top}$; this follows because $\sigma_{t}$ is a strictly increasing function of $|X_{t-1}|$ for all $t$ and copulas are invariant under strictly increasing transformations of variables. For the same reason, it is also the copula density for the variables $(X_{1}^{2},X_{2}^{2})^{\top}$.
Figure S2: Left: copula density for copula of the absolute values for classical ARCH(1) copula with Gaussian innovations; see Section 2.4.1 for parameters. Right: AR(1)-ARCH(1) model with Gaussian innovations and parameters $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
Appendix S3 General proof of Proposition 3
We complete the proof of Proposition 3. The case $k=2$ has been proved so let $k\geqslant 3$. Suppose we write $s_{2}(x,s)=\sigma(x,s)$, $s_{3}(z_{1},x,s)=\sigma(z_{1},s_{2}(x,s))$ and then for $j=4,\ldots,k+1$ we define recursive functions $s_{j}(z_{j-2},\ldots,z_{1},x,s)=\sigma(z_{j-2},s_{j-1}(z_{j-3},\ldots,z_{1},x,s))$. This permits us to write the joint density as
$f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)=\int_{0}^{\infty}f_{X_{k+1}\mid\sigma_{k+1}}(y\mid s_{k+1}(z_{k-1},\ldots,z_{1},x,s))\\ \left(\prod_{j=3}^{k}f_{X_{j}\mid\sigma_{j}}(z_{j-1}\mid s_{j}(z_{j-2},\ldots,z_{1},x,s))\right)f_{X_{2}\mid\sigma_{2}}(z_{1}\mid s_{2}(x,s))f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s.$
This satisfies $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(x,\ldots,-y)$ when the innovation distribution is symmetric. Moreover, in the symmetric model, we have $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(-x,\ldots,y)$ since all of the functions $s_{2},\ldots,s_{k+1}$ are even functions of $x$. These properties also hold for the conditional density
$f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}(x,y\mid z_{1},\ldots,z_{k-1})=f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)/f_{X_{2},\ldots,X_{k}}(z_{1},\ldots,z_{k-1}).$
To infer that the assertions regarding symmetry of the conditional copula hold we note that h-symmetry of $f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}$ implies that $F_{X_{k+1}\mid\bm{X}_{[2:k]}}$ is symmetric and joint symmetry of $f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}$ also implies that $F_{X_{1}\mid\bm{X}_{[2:k]}}$ is symmetric; the result then follows easily from equation (10).
Since the joint density of $(X_{1},X_{k+1})^{\top}$ can be obtained as
it follows by similar logic and application of Lemma 1 that the copula $C_{(k)}$ of the unconditional distribution of $(X_{1},X_{k+1})^{\top}$ shares the same symmetry properties as the conditional copula $C_{k}$.
Appendix S4 The absolute spherical t copula
The absolute spherical t copula has density given by
@@ 5 Conclusion
It has an asymptote at $(1,1)$ but takes a finite value at $(0,0)$ given by
$c^{|t|}_{\nu}(0,0)=\frac{\Gamma\left(\frac{\nu+2}{\nu}\right)\Gamma\left(\frac{\nu}{2}\right)}{\Gamma\left(\frac{\nu+1}{2}\right)^{2}}.$
This behaviour facilitates the maximum likelihood estimation of fulcrum parameters in models that incorporate v-transforms.
@@
Integrating any of the four equalities for $c^{|t|}_{\nu}$ and observing the boundary conditions for a copula gives us an expression for $C^{|t|}_{\nu}$. For example, integrating the final equality gives
$C^{|t|}_{\nu}(u,v)=4C_{\nu}^{t}\left(\frac{1+u}{2},\frac{1+v}{2}\right)-u-v-1.$ (S1)
The h-functions for the copula may be calculated to be
$h^{|t|}_{\nu,1}(u,v)$ $=\frac{\partial}{\partial u}C^{|t|}_{\nu}(u,v)=2t_{\nu+1}\left(\sqrt{\frac{\nu+1}{\nu+t_{\nu}^{-1}(\frac{1+u}{2})}}t_{\nu}^{-1}\left(\frac{1+v}{2}\right)\right)-1$ (S2)
$h^{|t|}_{\nu,2}(u,v)$ $=\frac{\partial}{\partial v}C^{|t|}_{\nu}(u,v)=2t_{\nu+1}\left(\sqrt{\frac{\nu+1}{\nu+t_{\nu}^{-1}(\frac{1+v}{2})}}t_{\nu}^{-1}\left(\frac{1+u}{2}\right)\right)-1$
where $t_{\nu}$ denotes the distribution function of a Student t distribution with $\nu$ degrees of freedom.
For calculation purposes, the h-functions are easy to evaluate for all $\nu>0$, since the function $t_{\nu}$ is typically available in statistical software. This is not however the case for the copula $C^{|t|}_{\nu}$ itself, since available software for evaluating the t copula $C^{t}_{\nu,\rho}$ is currently restricted. The R package copula is restricted to integer $\nu$ while the package rvinecopulib is confined to $\nu\geqslant 2$. A one-dimensional numerical integral seems unavoidable for evaluating the copula for general $\nu$ and we solve this problem by integrating the h-functions in (S2).
From expression (S1) we can infer that the absolute t copula interpolates between independence (as $\nu\to\infty$) and comonotonicity (as $\nu\to 0$). For the latter case this follows because the spherical t copula $C_{\nu}^{t}$ converges to $0.5M(u,v)+0.5W(u,v)$, an equal mixture of the comonotonicity and countermonotonicity copulas, as $\nu\to 0$ (McNeil et al.,, 2022).
In view of the limiting behaviour of the copula, the Kendall’s tau value $\tau(\nu)$ of $C^{|t|}_{\nu}$ must range in the interval $(0,1)$ as $\nu$ ranges in the interval $(0,\infty)$. We can find values by numerical integration using the formula
$\tau(\nu)=4\int_{0}^{1}\int_{0}^{1}C^{|t|}_{\nu}(u,v)c^{|t|}_{\nu}(u,v)\mathrm{d}u\mathrm{d}v-1=64\int_{0}^{0.5}\int_{0}^{0.5}C_{\nu}^{t}(u,v)c_{\nu}^{t}(u,v)\mathrm{d}u,\mathrm{d}v-1.$ (S3)
For example, for $\nu\in\{4,2,1,0.5\}$ the corresponding values are $\{0.099,0.189,0.333,0.515\}$. The absolute spherical t copula has upper tail dependence (inherited from the t copula). However we can easily show that it is asymptotically independent in the lower tail.
Appendix S5 Simulation study of the approximation of $C_{1}$
In the simulation study, data are generated from ARCH(1) and GARCH(1,1)-type processes using the parameter values summarized in Table 1 and different innovation distributions, as recorded in the legend to the table. Analyses are based on simulated samples of 50000 points and the first-order copula $C_{1}$ is estimated by fitting a first-order D-vine to the data using the two-stage method developed by Chen and Fan, (2006) in which the marginal distribution of the process is estimated by the rescaled empirical distribution function. Models are compared according to their AIC values. The results of the jointly symmetric case are documented in Section 3.5.1. Further experiments are documented here.
S5.1 Simulation study of h-symmetric case with leverage